OpenStudy (anonymous):

This year's soybean crop of 224, 000 tons represents an increase of 40% over last year's crop. How many tons of soybean were produce last year?

OpenStudy (anonymous):

well we take last years crop to be x tons, we know that this years crop was 140% of last years or 100% + 40%, we can represent 140% as a decimal by 1.40 so then using a we say x tons *1.40=224,000 tons then we divide by 1.4 and get x= 224000/1.4 which yields x=160,00 tons we can check this by saying 160000+(160000*0.4)=? 224000 so 160000*.4 = 64000 and 160000 + 64000 = 224000, so our answer checks